OpenStudy (anonymous):

3(x - 6) + 2 = 4(x + 2) - 21 x =

OpenStudy (anonymous):

parenthesis first. 3(x-6)= 3x-18 and 4(x+2)= 4x+8

OpenStudy (anonymous):

3x-18+2=4x+8-21

OpenStudy (anonymous):

combine like terms. 3x-16=4x-17

OpenStudy (anonymous):

3x-16=4x-17 -3x -3x -16=x-17 +16 +16 x= -1